the measures of 2 angles in a triangle are equal. The third angle measures 2 times the sum of the equal angles. Find the measures of the three angles.
Answer by mananth(16946)

the measures of 2 angles in a triangle are equal. The third angle measures 2 times the sum of the equal angles. Find the measures of the three angles.
Let the equal angles be x
Sum of two angles =2x
Third angle = 2*2x =4x
4x+2x=180 ( sum of all three angles of a triangle)
6x=180
X=30
Third angle = 120
30,30,120 deg